# ZS4C: Zero-Shot Synthesis of Compilable Code for Incomplete Code Snippets using ChatGPT

### v1

@@ 5.1. Results of RQ1
Table 2. The comparison of SnR, BasePrompt, and ZS4C in terms of F1, precision (Pre), and recall (Rec), and compilation rate (CR). We present the average for F1, Rec, Pre, and the total compilation rate in the summary row. The cells that have the best performance are marked in bold.
SnR BasePrompt ZS4C
@@
Android
0.952
0.909
1
37
0.967
0.95
0.987
44
0.990
0.990
0.990
48
JDK
0.887
0.796
1
12
0.482
0.317
1
15
0.997
1.000
0.994
23
Hibernate
0.948
0.918
0.980
23
0.75
0.606
0.985
22
0.902
0.834
0.982
32
Joda-Time
0.860
0.759
0.992
37
0.979
0.971
0.988
44
0.985
0.994
0.977
50
GWT
0.835
0.721
0.991
30
0.951
0.927
0.977
32
0.968
0.940
0.997
46
XStream
0.934
0.877
1
29
0.863
0.78
0.967
22
0.934
0.890
0.980
35
Summary
0.903
0.830
0.994
168 (62.9%)
0.832
0.759
0.984
179 (67%)
0.963 0.938
0.987
234 (87.6%)
[…]
Same Different Extra Missing None
SnR 143 78 4 35 7
BasePrompt 173 55 14 19 6
ZS4C 202 35 16 14 0
ImportStInfer 191 48 14 12 2
[…]
ConversationFixing 175 48 22 19 3
We present the comparison results among ZS4C and two baseline approaches in Tables 2 and  3.
The experimental results show that ZS4C successfully synthesized 234 compilable code snippets out of the 267 provided incomplete code snippets, resulting in an impressive overall compilation rate of 87.6%, outperforming two baselines. Compared with SnR, ZS4C showed a 39.3% improvement in terms of compilation rate. [^5] SnR paper reports that they can compile 197 out of 267 (73.8%) code snippets successfully after the repair. However, in their replication package, we found 168 code snippets were marked as comparable. Nevertheless, our approach outperforms SnR by improving the number of compilable code snippets to 234. Furthermore, ZS4C consistently outperformed the two approaches for all libraries. A major factor behind this boosting in compilation rate is the ConversationFixing component, which enables conversations between the compiler and the LLM. A deeper dive into the effectiveness and impact of this component can be found in RQ2 and Section 6.1.
When measuring the accuracy of the synthesized import statements, on average, ZS4C achieves 0.961, 0.987, and 0.938 in terms of F1, Precision, and Recall, respectively, as detailed in Table 2. Compared with SnR, ZS4C achieves an improvement of 6.6% and 13.4% in terms of F1 and precision. In terms of Recall, ZS4C has a slightly lower performance than SnR, which achieves an almost perfect precision of 0.994 due to its constraint-based inference. In six of the target libraries, ZS4C outperforms SnR in five, except Hibernate. We perform failure cases analysis to understand why eventually some cases are not compilable after the fixing (see more details in Section 6.1).
Compared with BasePrompt, ZS4C can compile 55 more code snippets, reflecting a 30.7% enhancement in the compilation rate. Furthermore, ZS4C elevates the accuracy of import statement inference, with the F1 score rising from 0.832 to 0.961. These improvements indicate the significance of task-specific prompt engineering in optimizing the performance of LLMs like ChatGPT.
To complement accuracy metrics, we also compare the list-wise matching degree. ZS4C (202) can infer more exact match import statement lists (same) than SnR (143) [^6] SnR paper reports that they can achieve 183 same, 62 different, 11 extra, 4 missing, and 7 none matches. However, in their replication package, we found the number we reported in the table. Nevertheless ZS4C can achieve more exact same matches. and BasePrompt (193). Also ZS4C can always output thing that is related to the correct import statements (none is 0).
To understand why ZS4C outperforms SnR, we manually study the cases where ZS4C performs better than SnR. We observe two reasons. First, SnR needs to repair an input code snippet to a syntactically valid compilation unit and then generate the Abstract Syntax Tree (AST) for the following procedures. However, the syntax repairing is not 100% guaranteed. Therefore, in certain cases, SnR failed in the first repairing step. While ZS4C does not rely on static analysis (e.g., parsing AST). LLM directly takes code snippets as sequence input and does the inference, which reduces the risk of failures due to errors incurring in static analysis. Second, SnR infers the appreciated import statements by solving constraints. However, if multiple libraries meet the constraints simultaneously, SnR does not have a strategy to handle this. While ZS4C leverages LLMs as the knowledge base to perform the inference, LLMs have sufficient knowledge to decide which libraries are most appreciated given the context.
ZS4C improves the compilation rate of the SOTA approach, SnR, from 63% to 87.6%. On average, ZS4C’s synthesized import statements are more accurate than SnR, with an improvement of 6.6% in the F1 score. Furthermore, ZS4C is more effective than BasePrompt, which represents a straightforward way of utilizing ChatGPT for this task.
5.2. RQ2: What is the contribution of individual components to ZS4C?
[…]
F1 Rec Pre CR F1 Rec Pre CR F1 Rec Pre CR
Android
0.948
0.915
0.982
39
0.980
0.980
0.980
42
0.927
0.902
0.954
43
JDK
0.919
0.850
1
19
0.997
1
0.994
23
0.825
0.707
0.992
21
Hibernate
0.886
0.813
0.974
27
0.898
0.828
0.981
38
0.876
0.803
0.963
34
Joda-Time
0.964
0.941
0.988
42
0.967
0.959
0.976
29
0.968
0.965
0.970
49
GWT
0.939
0.905
0.976
35
0.956
0.924
0.990
45
0.888
0.879
0.896
44
XStream
0.884
0.936
0.837
31
0.910
0.850
0.980
28
0.898
0.850
0.951
34
Summary
0.923
0.893
0.960
193 (72.2%)
0.951
0.924
0.984
205 (76.8%)
0.897
0.851
0.954
225 (84.2%)
[…]
We additionally observe that ConversationFixing achieves a higher compilation rate (84.2%) than ImportStInfer (72.2%). This underscores the benefits derived from the collaborative conversations between the LLM and the compiler. Indeed, ConversationFixing simulates a practical scenario where a developer interacts with ChatGPT through several rounds to address the compilation errors in their code. Yet, when we shift the focus to the efficacy of both components in inferring missing import statements, the tables turn. ImportStInfer has a F1 of 0.951, which is much higher than ConversationFixing’s F1 score of 0.897. This demonstrates that extracting all required important statements from the code body can be difficult when the primary goal of the LLM is to resolve compilation errors. By breaking the task into two steps—first inferring import statements and then iteratively fixing compilation errors—the efficiency of the LLM is enhanced. As shown in Table 4, when isolated, neither component approaches the performance of ZS4C in terms of the two considered criteria. This shows that both components significantly contribute to ZS4C, and their combined strengths boost ZS4C’s performance.
@@ 5.2. RQ2: What is the contribution of individual components to ZS4C?
5.3. RQ3: Does few-shot learning improve ZS4C?
Table 5 presents the results of ZS4C in two settings, zero-shot and one-shot. Surprisingly, one-shot in-context learning does not outperform zero-shot in any evaluation metrics. Instead, there is a slight decline in terms of F1. We suspect the reason could have originated from the advantages of using demonstration examples. Demonstration examples are usually provided for in-context learning for two purposes: 1) to help LLM understand the intent of a task and learn the patterns to complete the task (White et al., 2023; Webson and Pavlick, 2021; Lu et al., 2021); 2) to provide domain-specific knowledge for downstream tasks (Nashid et al., 2023; Gao et al., 2023; Zhou et al., 2022a). In our cases, apparently, none of the above advantages of providing examples hold. Our task is straightforward, and the model can understand the intent expressed in the zero-shot prompt precisely. This is possible because ChatGPT is trained on a large dataset Common Crawl, which includes data from Stack Overflow and Github (Nagel, 2022). ChatGPT could understand the task precisely. Meanwhile, the demonstrations do not reveal any domain-specific knowledge that can be utilized.
Table 5. The comparison of the performance of ZS4C in one-shot and zero-shot settings.
Compilation Rate F1 Recall Precision
one-shot 86.9% 0.955 0.934 0.981
zero-shot 87.6% 0.961 0.938 0.987
Surprisingly, one-shot in-context learning does not improve the performance of ZS4C.
5.4. RQ4: How do parameters affect the effectiveness of ZS4C?
Figure 3. The impact of parameters $K$ (left) and $M$ (right) on ZS4C.
Figure 3 (left) presents the results of parameter sensitivity analysis on $K$ (i.e., number of sampling in self-consistency). As $K$ increases, the effectiveness of ZS4C to infer the import statements declines slightly, from 0.961 to 0.934 in terms of F1. We can observe similar trends for accuracy, recall, and precision. However, we observe the compilation rate improves slightly from 0.873 to 0.891. One possible explanation is that ConversationFixing provides a safe net for ImportStInfer even if it makes mistakes for some cases.

### v3

@@ 5.1. RQ1: How effective is our approach ZS4C?
Table 3. The comparison of SnR, BasePrompt, ZS4C with GPT-3.5 and ZS4C with GPT-4 for StatType-SO (Java) dataset in terms of F1, recall (Rec), precision (Pre), and compilation rate (CR). We present the average for F1, Rec, Pre, and the total compilation rate in the summary row. The cells that have the best performance are marked in bold.
SnR BasePromptGPT-3.5 BasePromptGPT-4 ZS4CGPT-3.5 ZS4CGPT-4
F1 Rec Pre CR F1 Rec Pre CR F1 Rec Pre CR F1 Rec Pre CR F1 Rec Pre CR
Android 0.952 0.909 1 37 (74%) 0.967 0.95 0.987 44 (88%) 0.969 0.956 0.983 45 (90%) 0.990 0.990 0.990 48 (96%) 0.988 0.99 0.987 48 (96%)
JDK 0.887 0.796 1 12 (52.2%) 0.482 0.317 1 15 (65.2%) 0.713 0.557 0.989 18 (78.3%) 0.997 1.000 0.994 23 (100%) 0.994 1 0.988 23 (100%)
Hibernate 0.948 0.918 0.98 23 (46%) 0.75 0.606 0.985 22 (44%) 0.834 0.747 0.944 29 (58%) 0.902 0.834 0.982 32 (64%) 0.956 0.916 1 41 (82%)
Joda-Time 0.86 0.759 0.992 37 (74%) 0.979 0.971 0.988 44 (88%) 0.991 0.994 0.988 49 (98%) 0.985 0.994 0.977 50 (100%) 0.994 0.988 1 50 (100%)
GWT 0.835 0.721 0.991 30 (60%) 0.951 0.927 0.977 32 (64%) 0.906 0.889 0.924 42 (84%) 0.968 0.94 0.997 46 (92%) 0.987 0.978 0.997 50 (100%)
XStream 0.934 0.877 1 29 (65.9%) 0.863 0.78 0.967 22 (50%) 0.927 0.89 0.967 38 (86.4%) 0.934 0.89 0.98 35 (79.6%) 0.959 0.925 0.995 42 (95.5%)
Summary 0.903 0.83 0.994 168 (62.9%) 0.832 0.759 0.984 179 (67%) 0.890 0.839 0.966 221 (82.8%) 0.963 0.938 0.987 234 (87.6%) 0.98 0.966 0.995 254 (95.1%)
Table 4. The comparison of BasePrompt (with GPT-3.5 and GPT-4) and ZS4C (with GPT-3.5 and GPT-4) for Python dataset (Python-SO) in terms of F1, recall (Rec), precision (Pre), and compilation rate (CR). We present the average for F1, Rec, Pre, and the total compilation rate in the summary row. The cells that have the best performance are marked in bold.
BasePromptGPT-3.5 BasePromptGPT-4 ZS4CGPT-3.5 ZS4CGPT-4
F1 Rec Pre CR F1 Rec Pre CR F1 Rec Pre CR F1 Rec Pre CR
BeautifulSoup 0.954 0.913 1 29 (82.9%) 0.947 0.9 1 31 (88.6%) 0.968 0.938 1 31 (88.6%) 0.974 0.95 1 33 (94.3%)
Flask 0.937 0.908 0.967 15 (65.2%) 0.977 0.985 0.97 18 (78.3%) 1 1 1 23 (100%) 1 1 1 23 (100%)
Gensim 0.919 0.895 0.944 10 (90.9%) 0.842 0.842 0.842 9 (81.8%) 1 1 1 11 (100%) 0.973 0.947 1 11 (100%)
Keras 0.947 0.953 0.942 15 (88.2%) 0.948 0.965 0.932 13 (76.5%) 0.988 0.988 0.988 14 (82.4%) 0.994 0.988 1 15 (88.2%)
LightGBM 0.947 0.9 1 6 (85.7%) 0.976 1 0.952 6 (85.7%) 1 1 1 7 (100%) 1 1 1 7 (100%)
Matplotlib 0.983 0.978 0.989 49 (100%) 0.989 0.989 0.989 49 (100%) 0.994 0.989 1 46 (93.9%) 0.994 0.989 1 47 (95.9%)
NLTK 0.968 0.957 0.978 10 (55.6%) 0.957 0.957 0.957 13 (72.2%) 1 1 1 18 (100%) 0.978 0.957 1 18 (100%)
Numpy 0.887 0.796 1 15 (88.2%) 0.887 0.796 1 17 (100%) 0.826 0.704 1 16 (94.1%) 0.839 0.722 1 17 (100%)
OpenCV 0.984 0.989 0.978 25 (50%) 0.989 1 0.979 31 (62%) 1 1 1 48 (96%) 1 1 1 49 (98%)
Pandas 0.973 0.947 1 7 (87.5%) 0.947 0.947 0.947 8 (100%) 0.973 0.947 1 7 (87.5%) 0.973 0.947 1 7 (87.5%)
Plotly 0.92 0.886 0.956 24 (45.3%) 0.946 0.919 0.974 27 (50.9%) 0.94 0.887 1 40 (75.5%) 0.937 0.896 0.982 42 (79.2%)
Pytorch 0.945 0.968 0.923 23 (88.5%) 0.903 0.903 0.903 25 (96.2%) 0.984 0.968 1 26 (100%) 0.949 0.903 1 26 (100%)
Scikit-learn 1 1 1 17 (100%) 1 1 1 17 (100%) 1 1 1 17 (100%) 0.989 0.979 1 17 (100%)
Scipy 0.963 0.948 0.979 24 (64.9%) 0.958 0.938 0.978 26 (70.3%) 0.974 0.949 1 37 (100%) 0.979 0.959 1 37 (100%)
Scrapy 0.875 0.875 0.875 32 (65.3%) 0.877 0.892 0.863 34 (69.4%) 0.966 0.934 1 47 (95.9%) 0.97 0.942 1 49 (100%)
Seaborn 0.974 0.974 0.974 45 (93.7%) 0.965 0.974 0.957 47 (97.9%) 0.987 0.974 1 48 (100%) 0.987 0.974 1 48 (100%)
Spacy 0.926 0.875 0.982 18 (69.2%) 0.935 0.906 0.967 25 (96.2%) 0.984 0.969 1 16 (61.5%) 0.984 0.969 1 23 (88.5%)
Tensorflow 0.961 0.961 0.961 7 (43.8%) 0.949 0.922 0.979 10 (62.5%) 0.99 0.98 1 13 (81.3%) 1 1 1 15 (93.8%)
Theano 0.972 0.946 1 9 (81.8%) 0.946 0.946 0.946 8 (72.7%) 0.974 0.974 1 8 (72.7%) 0.986 0.973 1 10 (90.9%)
XGBboost 0.971 0.944 1 17 (81%) 0.965 0.958 0.971 18 (85.7%) 0.986 0.972 1 18 (85.7%) 0.986 0.972 1 20 (95.2%)
Summary 0.95 0.931 0.972 397 (73.7%) 0.945 0.937 0.955 432 (80.2%) 0.977 0.959 0.999 491 (91.1%) 0.975 0.953 0.999 514 (95.4%)
Table 5. The comparison of SnR, BasePrompt, ZS4C in terms of list-wise matching degree on StatType-SO and Python-SO datasets.
StatType-SO
Approach Same Different Extra Missing None
SnR 143 (53.56%) 78 (29.21%) 4 (1.5%) 35 (13.11%) 7 (2.62%)
BasePromptGPT-3.5 173 (64.79%) 55 (20.60%) 14 (5.24%) 19 (7.12%) 6 (2.25%)
BasePromptGPT-4 226 (84.64%) 24 (8.99%) 10 (3.75%) 7 (2.62%) 0 (0%)
ZS4CGPT-3.5 202 (75.66%) 35 (13.11%) 16 (5.99%) 14 (5.24%) 0 (0%)
ZS4CGPT-4 226 (84.64%) 24 (8.99%) 10 (3.75%) 7 (2.62%) 0 (0%)
Python-SO
Approach Same Different Extra Missing None
BasePromptGPT-3.5 447 (82.93%) 43 (7.98%) 27 (5.01%) 19 (3.53%) 3 (0.56%)
BasePromptGPT-4 447 (82.93%) 42 (7.79%) 39 (7.24%) 9 (1.67%) 2 (0.37%)
ZS4CGPT-3.5 491 (91.09%) 35 (6.49%) 4 (0.74%) 9 (1.67%) 0 (0%)
ZS4CGPT-4 499 (92.58%) 30 (5.57%) 3 (0.56%) 7 (1.30%) 0 (0%)
We present the comparison results for StatType-SO in Tables 3 and for Python-SO in Table 4. The results show that both ZS4CGPT-3.5 and ZS4CGPT-4 outperform the baseline SnR by a large margin on StatType-SO. For instance, ZS4CGPT-3.5 successfully synthesized 234 compilable code snippets out of the 267 provided incomplete code snippets, resulting in an impressive overall compilation rate of 87.6%, outperforming two baselines. Compared with SnR, ZS4CGPT-3.5 showed a 39.3% improvement in terms of compilation rate. [^7] SnR paper reports that they can compile 197 out of 267 (73.8%) code snippets successfully after the repair. However, in their replication package, we found 168 code snippets were marked as comparable. Nevertheless, our approach outperforms SnR by improving the number of compilable code snippets to 234. Furthermore, ZS4CGPT-3.5 consistently outperformed the two approaches for all libraries. ZS4CGPT-4 achieves a compilation performance improvement of 8.56% over its predecessor GPT-3.5, increasing from 87.6% to 95.1%. Similarly, on Python-SO, ZS4C outperforms BasePrompt. For instance, when using GPT3.5 as the base model, ZS4C achieves a compilation rate of 91.1%
When measuring the accuracy of the synthesized import statements, ZS4C always outperforms baselines in all evaluation metrics. For instance, on average, ZS4CGPT-4 achieves 0.98, 0.966, and 0.995 in terms of F1, Precision, and Recall on StatType-SO, respectively, as detailed in Table 3. Compared with SnR, ZS4CGPT-4 achieves an improvement of 8.5% and 16.4% in terms of F1 and recall. ZS4CGPT-4 and SnR have similar precision. In six of the target libraries, ZS4CGPT-4 outperforms SnR in all six libraries. We perform failure cases analysis to understand why eventually some cases are not compilable after the fixing (see more details in Section 6.2).
Compared with BasePrompt, ZS4C consistently can predict import statements more accurately and compile more code snippets. For instance, on StatType-SO ZS4CGPT-3.5 can compile 55 more code snippets, reflecting a 30.7% enhancement in the compilation rate, compared with BasePromptGPT-3.5. Furthermore, ZS4CGPT-3.5 elevates the accuracy of import statement inference, with the F1 score rising from 0.832 to 0.961. These improvements indicate the significance of task-specific prompt engineering in optimizing the performance of LLMs. We observe a similar trend on Python-SO dataset. ZS4C consistently outperforms BasePrompt with the same base model.
To complement accuracy metrics, we also compare the list-wise matching degree and present the result in Table 5. Both ZS4CGPT-4 (226) and ZS4CGPT-3.5 (202) can infer more exact match import statement lists (same) than SnR  [^8] SnR paper reports that they can achieve 183 same, 62 different, 11 extra, 4 missing, and 7 none matches. However, in their replication package, we found the number we reported in the table. Nevertheless ZS4C with GPT-3.5 can achieve more exact same matches. and BasePrompt (193) on StatType-SO. We observe a similar trend on Python-SO, ZS4C always outperforms BasePrompt. Also ZS4C can always output things that are related to the correct import statements (none is 0).
To understand why ZS4C outperforms SnR, we manually study the cases where ZS4C performs better than SnR. Since ZS4CGPT-4 and ZS4CGPT-3.5 achieve similar performance, here we study the cases from ZS4CGPT-3.5. We observe two reasons. First, SnR needs to repair an input code snippet to a syntactically valid compilation unit and then generate the Abstract Syntax Tree (AST) for the following procedures. However, the syntax repairing is not 100% guaranteed. Therefore, in certain cases, SnR failed in the first repairing step. While ZS4C with GPT-3.5 does not rely on static analysis (e.g., parsing AST). LLM directly takes code snippets as sequence input and does the inference, which reduces the risk of failures due to errors incurring in static analysis. Second, SnR infers the appreciated import statements by solving constraints. However, if multiple libraries meet the constraints simultaneously, SnR does not have a strategy to handle this. While ZS4C leverages LLMs as the knowledge base to perform the inference, LLMs have sufficient knowledge to decide which libraries are most appreciated given the context.
In addition, we observe that ZS4CGPT-4 consistently outperforms ZS4CGPT-3.5 or achieves comparable effectiveness as ZS4CGPT-3.5 on both studied datasets, typically in terms of compilation rate. ZS4CGPT-4 improves the compilation rate of ZS4CGPT-3.5 from 87.6% to 95.1%. This probably originates from two advantages of GPT-4. First, GPT-4 is more advanced than GPT-3.5 and has already proved more effective in various code-related tasks (Liu et al., 2024b; Poldrack et al., 2023). Second, GPT-4 has a long window size and can fit much more context in the prompt (e.g., history generated during Conversational Error Fixing).
In general, ZS4C with GPT-4 outperforms GPT-3.5 and other baselines on both StatType-SO and Python-SO. For instance, on StatType-SO, ZS4C with GPT-4 improves the compilation rate of the SOTA approach, SnR, from 63% to 95.1%. On average, ZS4C’s (with GPT-4) synthesized import statements are more accurate (with an F1 score of 0.98) than SnR, with an improvement of 8.5% in the F1 score. Furthermore, ZS4C is more effective than BasePrompt.
5.2. RQ2: What is the contribution of individual components to ZS4C?
Table 6. The comparison of ImportStInfernoSC, ImportStInfer, ConversationFixing in terms of F1, recall (Rec), precision (Pre), and compilation rate (CR) with GPT-3.5. We present the average for F1, Rec, and Pre and the total compilation rate in the summary row.
ImportStInfernoSC ImportStInfer ConversationFixing
@@ 5.2. RQ2: What is the contribution of individual components to ZS4C?
StatType-SO 0.923 0.893 0.960 193 (72.2%) 0.951 0.924 0.984 205 (76.8%) 0.897 0.851 0.954 225 (84.2%)
Python-SO 0.956 0.937 0.978 417 (77.4%) 0.965 0.942 0.993 451 (83.7%) 0.942 0.908 0.983 463 (85.9%)
Table 6 presents the results of the ablation analysis of ZS4C with GPT-3.5 on StatType-SO and Python-SO datasets. The experimental results show that the self-consistency method can improve the effectiveness of the Import Statement Inference component. Specifically, ImportStInfer has a 6.4% and 2% higher of compilation rate than ImportStInfernoSC on StatType-SO and Python-SO, respectively. The self-consistency method improves the results of ImportStInfernoSC from 0.923 to 0.951 and 0.956 to 0.965 in terms of average F1, on StatType-SO and Python-SO, respectively.
We additionally observe that ConversationFixing achieves a higher compilation rate than ImportStInfer on both datasets. This underscores the benefits derived from the collaborative conversations between the LLM and the compiler. Indeed, ConversationFixing simulates a practical scenario where a developer interacts with ChatGPT through several rounds to address the compilation errors in their code. Yet, when we shift the focus to the efficacy of both components in inferring missing import statements, the tables turn. On both datasets, ImportStInfer has a higher F1 score than ConversationFixing. This demonstrates that extracting all required important statements from the code body can be difficult when the primary goal of the LLM is to resolve compilation errors. By breaking the task into two steps—first inferring import statements and then iteratively fixing compilation errors—the efficiency of the LLM is enhanced. As shown in Table 6, when isolated, neither component approaches the performance of ZS4C in terms of the two considered criteria. This shows that both components significantly contribute to ZS4C, and their combined strengths boost ZS4C’s performance.
@@
5.3. RQ3: How do parameters affect the effectiveness of ZS4C?
Figure 3. The impact of parameters $K$ (left) and $M$ (right) on ZS4CGPT-3.5 for StatType-SO.
Figure 3 (left) presents the results of parameter sensitivity analysis on $K$ (i.e., number of sampling in self-consistency). As $K$ increases, the effectiveness of ZS4C with GPT-3.5 for StatType-SO (Java) dataset to infer the import statements declines slightly, from 0.961 to 0.934 in terms of F1. We can observe similar trends for accuracy, recall, and precision. However, we observe the compilation rate improves slightly from 0.873 to 0.891. One possible explanation is that ConversationFixing provides a safe net for ImportStInfer even if it makes mistakes for some cases.
@@ 5.3. RQ3: How do parameters affect the effectiveness of ZS4C?
5.4. RQ4: Does few-shot learning improve ZS4C?
Table 8 presents the results of ZS4C in three settings, zero-shot, one-shot, and three-shot (few-shot) of ZS4CGPT-3.5 and ZS4CGPT-4 on both datasets.
Surprisingly, one-shot in-context learning does not outperform zero-shot with GPT-3.5 in any evaluation metrics. Instead, there is a slight decline in terms of F1. With GPT-4, we observe a similar trend, where few-shot in-context learning does not improve the performance of ZS4C. Instead, the performance declines in the three-shot setting compared with zero-shot in terms of CR on StatType-SO. One possibility is that the code retrieved is not similar to the input code so the few-shot setting does not help. To verify this, we examine the similarity between each input code and their corresponding retrieved code snippet. Table 7 presents the similarity score and we observe that the retrieved code is actually similar to the input code snippet. So this assumption is rejected. Then, we suspect the reason could be that the demonstration examples do not provide any additional information to the model. Demonstration examples are usually provided for in-context learning for two purposes: 1) to help LLM understand the intent of a task and learn the patterns to complete the task (White et al., 2023; Webson and Pavlick, 2021; Lu et al., 2021); 2) to provide domain-specific knowledge for downstream tasks (Nashid et al., 2023; Gao et al., 2023; Zhou et al., 2022a). In our cases, apparently, none of the above advantages of providing examples hold. The model can understand the intent expressed in the zero-shot prompt accurately because of the clear problem definition. The reason may be that GPT is large enough to understand many programming tasks precisely. We demonstrate that with a powerful model, zero-shot synthesis of compilable code for incomplete code snippets using LLMs is feasible.
Table 7. Summary of similarity score between each input code and their corresponding retrieved code snippet for one-shot setting in StatType-SO and Python-SO datasets.
StatType-SO
Library Minimum Mean Maximum
Android 0.901 0.962 0.985
JDK 0.791 0.924 0.973
Hibernate 0.912 0.95 0.984
JodaTime 0.937 0.963 0.986
GWT 0.935 0.965 0.984
XStream 0.938 0.953 0.983
Python-SO
Library Minimum Mean Maximum
BeautifulSoup 0.912 0.953 0.98
Flask 0.914 0.952 0.984
Gensim 0.904 0.931 0.945
Keras 0.867 0.945 0.985
Lightgbm 0.922 0.944 0.958
Matplotlib 0.816 0.932 0.989
NLTK 0.876 0.941 0.972
Numpy 0.85 0.932 0.968
OpenCV 0.81 0.954 0.986
Pandas 0.887 0.951 0.988
Plotly 0.928 0.964 0.986
Pytorch 0.912 0.937 0.956
Scikit-learn 0.894 0.94 0.986
Scipy 0.845 0.937 0.981
Scrapy 0.869 0.957 0.983
Seaborn 0.727 0.949 0.988
Spacy 0.695 0.936 0.97
Tensorflow 0.848 0.949 0.985
Theano 0.891 0.938 0.951
XGBoost 0.916 0.958 0.988
Table 8. The comparison of the performance of ZS4C in one-shot and zero-shot settings. Here, CR means compilation rate.
StatType-SO Python-SO
CR F1 Recall Precision CR F1 Recall Precision
Zero-shot (GPT-3.5) 87.6% 0.961 0.938 0.987 91.1% 0.977 0.959 0.999
One-shot (GPT-3.5) 86.9% 0.955 0.934 0.981 90.2% 0.958 0.942 0.979
Zero-shot (GPT-4) 95.1% 0.98 0.966 0.995 95.2% 0.975 0.953 0.999
One-shot (GPT-4) 95.9% 0.985 0.974 0.999 94.8% 0.976 0.96 0.995
Three-shot (GPT-4) 93.6% 0.986 0.974 0.999 95.4% 0.973 0.963 0.996
Surprisingly, for both models GPT-3.5 and GPT-4, few-shot in-context learning does not improve the performance of ZS4C. Typically, with GPT-4, ZS4C can achieve a high F1 and Compilation Rate with zero shot.